OpenStudy (anonymous):

Change the logarithmic expression to an equivalent expression involving an exponent. log3a9=2/3 Change the logarithmic expression to an equivalent expression involving an exponent. log3a9=2/3 @Mathematics

jhonyy9 (jhonyy9):

log3a9=2/3log3(3) log3a9 = log3(3)(2/3) a9=cuberoot9

OpenStudy (anonymous):

Thank you